A man is pulling a crate along a concrete floor. This time, let's be specific. The crate has a mass of 100 kg and the man pulls with a force of 1 250 N. The coefficient of friction between the crate and the floor is 0.2. What is the acceleration of the crate?


1
Expert's answer
2021-02-10T18:22:49-0500
"F_{pull}-F_{fr}=ma,""F_{pull}-\mu mg=ma,""a=\dfrac{F_{pull}-\mu mg}{m},""a=\dfrac{1250\ N-0.2\cdot100\ kg\cdot9.8\ \dfrac{m}{s^2}}{100\ kg}=10.54\ \dfrac{m}{s^2}."